How much energy is required to heat a 0.3kg package of frozen vegetable juice to 5 degrees celsius if it is initially at -10 degrees celsius. Assume that the vegetable juice is essentially water such that Cice = 2100 kg/Celsius, Lf= 334,000J/kg, and Cwatee = 4186 J/kg Celsius.
Answer #1

here,

mass of frozen vegetable juice , m = 0.3 kg

initial temperature , Ti = - 10 degree C

final temperature , Tf = 5 degree C

the energy required , E = m * ( Ci * ( 0 - Ti) + Lf + Cw * ( Tf - 0))

E = 0.3 * ( 2100 * 10 + 334000 + 4186 * 5) J

E = 1.13 * 10^5 J